d^2y/dx^2-6dy/dx+9y=1+x+x^2
1
Expert's answer
2021-09-19T18:07:35-0400

"\frac{d^{2}y}{dx^2}-6\frac{dy}{dx}+9y=1+x+x^{2}"

P(s)=s2-6s+9

=(s-3)2

Since we have two identical roots,the homogeneous solution will have the form ;

yh=C1e3x + C2xe3x

yp=Ax2+Bx+C

y'p=2Ax+B

y''p=2A

9Ax2+(9B-12A)x+(2A-6B+9C)=x2+x+1

9A=1 "\implies" A="\frac{1}{9}"

9B-12A=1 "\implies B=\frac{7}{27}"

2A-6B+9C=1"\implies C=\frac{7}{27}"

So yp="\frac{1}{9}x^{2}+\frac{7}{27}x+\frac{7}{27}"

"\therefore" The general solution is:

"y=C_1e^{3x}+C_2xe^{3x}+\frac{1}{9}x^{2}+\frac{7}{27}x+\frac{7}{27}"